OpenStudy (anonymous):

There are 3 red and 8 green balls in a bag. If you randomly choose balls one at a time, with replacement, what is the probability of choosing 2 red balls and then 1 green ball?

OpenStudy (anonymous):

multiply your probabilities together

OpenStudy (anonymous):

3/11 * 3/11 * 8/11 = 72/1331

OpenStudy (anonymous):

\[\frac{3}{11}\times \frac{3}{11}\times \frac{8}{11}\]

OpenStudy (anonymous):

Suppose that 4.5% of the people in the world have a particular genetic defect and that a screening test is 81% accurate for people who have it and 88% accurate for people who do not. If 1300 people are screened for the defect, which is the best prediction for the number of people without the defect who are identified as having it?

OpenStudy (kropot72):

Probability of having defect = 0.045 Probability of not having defect = (1 - 0.045) Predicted number of screened people who don't have defect = 1300(1 - 0.045) Predicted number of correct negative results = 1300(1 - 0.045) * 0.88 Predicted number of false positive results = 1300 - (1300 * 0.955 * 0.88) = ? Can you finish it?